OpenStudy (anonymous):

A video game company has determined that the revenue generated by a particular video game is modeled by the following equation: r(x)=10x-0.0025x^2 The variable x is measured in thousands of video games produced, and r(x) is measured in thousands of dollars. What is the maximum revenue the company can earn with this video game?

OpenStudy (anonymous):

use the formula x = -b / ( 2a) for y = ax^2 +bx + c . this is the maximum x value (in thousands of games) . to find the max revenue plug this baby into r(x). so x = -10 / ( 2*-.0025)

OpenStudy (anonymous):

Set the derivative of the expression to zero and solve for x. x = 2000 A plot is attached.

OpenStudy (anonymous):

how did you graph this ?

OpenStudy (anonymous):

Thanks, you were both very helpful!

OpenStudy (anonymous):

@ cantorset2 Used the Plot[ ] function from my copy of Mathematica 8, Home Edition.